# To CoT or not to CoT? Chain-of-thought helps mainly on math and symbolic reasoning

### v1

@@ Abstract
Chain-of-thought (CoT) via prompting is the de facto method for eliciting reasoning capabilities from large language models (LLMs). But for what kinds of tasks is this extra “thinking” really helpful? To analyze this, we conducted a quantitative meta-analysis covering over 100 papers using CoT and ran our own evaluations of 20 datasets across 14 models. Our results show that CoT gives strong performance benefits primarily on tasks involving math or logic, with much smaller gains on other types of tasks. On MMLU, directly generating the answer without CoT leads to almost identical accuracy as CoT unless the question or model’s response contains an equals sign, indicating symbolic operations and reasoning. Following this finding, we analyze the behavior of CoT on these problems by separating planning and execution and comparing against tool-augmented LLMs. Much of CoT’s gain comes from improving symbolic execution, but it underperforms relative to using a symbolic solver. Our results indicate that CoT can be applied selectively, maintaining performance while saving inference costs. Furthermore, they suggest a need to move beyond prompt-based CoT to new paradigms that better leverage intermediate computation across the whole range of LLM applications.
@@ 1 Introduction
How can we better understand why CoT improves on these questions and only these questions? The math and formal logical reasoning datasets we consider can be broken down into two stages of processing: a planning step (e.g., parsing a problem into equations) and an execution step (building intermediate outputs and working towards a solution) (Ye et al., 2023; Wang et al., 2023b; Sun et al., 2024). Finding 2: CoT primarily helps with the execution step that performs computation and symbolic manipulation, but falls short of what LLMs with tool augmentation can do. We find that LMs prompted with CoT can generate executable formal solution plans and execute those plans better than direct answering. But using LMs to generate a solution plan and then using an external symbolic solver to solve the plan outperforms using CoT for both stages for these tasks.
@@ 2 Background: Chain-of-thought
For the reasoning tasks we consider in this work, the output $\tilde{\mathbf{y}}$ can take one of multiple forms. A direct answer only contains a string realization of $a$; e.g., $\mathbf{y}=(\mathrm{\_185},\mathrm{4})$ which is detokenized as the answer $a=1854$. A chain of thought is a longer sequence $\mathbf{y}$ including other tokens beyond the answer, e.g., $\mathbf{y}=(\mathrm{\_185},\mathrm{6},\mathrm{\_minus},\mathrm{\_2},\mathrm{\_equals},\mathrm{\_185},\mathrm{4})$. [^2] In this paper, we only consider CoTs that end with an answer. Generating a prediction first followed by an explanation differs only minorly from direct answer in the few-shot setting (Ye & Durrett, 2022). In both cases, the $\mathrm{extract}$ function must detokenize the output and map the output to the correct datatype; in chain of thought, there is some extra work to spot where the answer is placed, though this can usually be done heuristically.
Our prompts can explicitly encourage use of direct answer or chain of thought as strategies, which we denote as $\mathcal{I}_{\mathrm{da}}$ and $\mathcal{I}_{\mathrm{cot}}$. For eliciting CoT, this includes strategies like telling a model to “think step by step(Kojima et al., 2022). For directly answering a question, a prompt may say “immediately generate the answer”. In this latter case, we form $\mathcal{I}_{\mathrm{da}}(\mathbf{q})=v_{\mathrm{da}}(\mathbf{q})$ using a verbalizer $v_{\mathrm{da}}$, which might concatenate the question with a description of the strategy like this (analogously for $\mathcal{I}_{\mathrm{cot}}$ and $v_{\mathrm{cot}}$). It is not guaranteed that LLMs prompted to use a certain strategy actually do. However, modern LLMs follow these instructions in the vast majority of cases. We track the average location of the answer in the generated output for both CoT and direct prompts in Appendix C.3 to ensure that direct answer prompts give the answer early in the output. We also ensure that $\mathrm{extract}$ can parse answers from the generated output for each model, prompt, and dataset combination used in our experiments, tailoring the $\mathrm{extract}$ function as needed to ensure low unparseable rates for each model and task. [^3] We exclude a number of other “CoT-like” approaches in our analysis such as decomposed prompting (Khot et al., 2023; Zheng et al., 2024a) and multi-agent debate (Du et al., 2023; Chen et al., 2024). We focus on single prompt approaches. We deal with tool-augmented approaches in Section 5. All prompts and outputs per dataset per model have been uploaded to Huggingface and we include examples of some of our prompts in the Appendix H. We also experiment with few-shot CoT prompts, which we find perform similarly to zero-shot prompts; details about these are given in Appendix B.
@@
Central claims
Figure 1 shows that there are a large number of positive results on CoT reported in the literature. Informally, we believe many readers of the literature to hold the following view:
Central Conjecture $\mathcal{I}_{\mathrm{cot}}$ will outperform $\mathcal{I}_{\mathrm{da}}$ on nearly all reasoning problems, whether those problems involve symbolic or non-symbolic reasoning.
Our evidence does not support this conjecture. We will show that this performance boost is strongest for symbolic and semi-symbolic tasks, while giving little to no improvement (or even hurting performance) on questions that are non-symbolic tasks.
@@ 3 Results from the Literature
We first perform a meta-analysis of recent papers comparing the performance of prompts $\mathcal{I}_{\mathrm{cot}}$ and $\mathcal{I}_{\mathrm{direct}}$ to identify the types of tasks where CoT has been reported to help.
3.1 Criteria and Process
@@
This resulted in a total of 1,218 experimental comparisons across 110 papers (35 from ICLR and 75 from NAACL and EACL) covering 264 datasets. Details and more information can be found in our GitHub Repo: https://github.com/Zayne-sprague/To-CoT-or-not-to-CoT.
Table 1: A few categories and their descriptions used to classify experimental comparisons. The complete list of categories and their descriptions is given in Appendix D.
@@
Logical reasoning Tasks designed to test for logical reasoning, whether deductive (e.g., PrOntoQA), inductive (Bowen et al., 2024) or analogical (Ma et al., 2024) reasoning, including syllogisms and logical puzzles.
@@
Categorization
Given the large number of tasks and datasets being compared, we grouped each task into a set of 14 categories. These categories were determined based on the description (and possibly examples) of the task, not taking into account system performance. These categories abstract over traditional NLP task classifications (e.g., NER, reading comprehension) and take into account both the task format and the kinds of reasoning involved. Definitions for several categories are shown in Table 1 and the full description is given in Appendix D.
3.2 Results
Do any non-math datasets benefit from CoT?
There are a few outliers that less clearly follow the trend. ScienceQA (Lu et al., 2022) consists of multiple choice questions across a range of natural and social science disciplines, though it is hard to interpret gains without knowing breaking down performance by subject or question type. The dialogue evaluation dataset from Jia et al. (2024) sees large improvements with CoT, but this is a proprietary dataset, and we note that other essay scoring results in our meta-analysis (Li et al., 2024; Stahl et al., 2024) did not show improvements with CoT. Other non-math, symbolic or logical datasets that benefit from CoT are Commitment Bank (de Marneffe et al., 2019) and the task of eliciting verbalized confidence (Xiong et al., 2024).
Nevertheless, these are exceptions to the rule. The majority of the reported benefits from using CoT in the NLP and ML literature comes from math or math-related tasks.

### v2

@@ Abstract
Chain-of-thought (CoT) via prompting is the de facto method for eliciting reasoning capabilities from large language models (LLMs). But for what kinds of tasks is this extra “thinking” really helpful? To analyze this, we conducted a quantitative meta-analysis covering over 100 papers using CoT and ran our own evaluations of 20 datasets across 14 models. Our results show that CoT gives strong performance benefits primarily on tasks involving math or logic, with much smaller gains on other types of tasks. On MMLU, directly generating the answer without CoT leads to almost identical accuracy as CoT unless the question or model’s response contains an equals sign, indicating symbolic operations and reasoning. Following this finding, we analyze the behavior of CoT on these problems by separating planning and execution and comparing against tool-augmented LLMs. Much of CoT’s gain comes from improving symbolic execution, but it underperforms relative to using a symbolic solver. Our results indicate that CoT can be applied selectively, maintaining performance while saving inference costs. Furthermore, they suggest a need to move beyond prompt-based CoT to new paradigms that better leverage intermediate computation across the whole range of LLM applications [^1] Our code can be found at https://github.com/Zayne-sprague/To-CoT-or-not-to-CoT..
@@ 1 Introduction
How can we better understand why CoT improves on these questions and only these questions? The math and formal logical reasoning datasets we consider can be broken down into two stages of processing: a planning step (e.g., parsing a problem into equations) and an execution step (building intermediate outputs and working towards a solution) (Ye et al., 2023; Wang et al., 2023b; Sun et al., 2024). Finding 2: CoT primarily helps with the execution step that performs computation and symbolic manipulation, but falls short of what LLMs with tool augmentation can do. We find that LMs prompted with CoT can generate executable formal solution plans and execute those plans better than direct answering. But using LMs to generate a solution plan and then using an external symbolic solver to solve the plan outperforms using CoT for both steps for these tasks.
@@ 2 Background: Chain-of-thought
For the tasks we consider in this work, the output $\tilde{\mathbf{y}}$ can take one of two forms. A direct answer only contains a string realization of $a$; e.g., $\mathbf{y}=(\mathrm{\_185},\mathrm{4})$ which is detokenized as the answer $a=1854$. A chain of thought is a longer sequence $\mathbf{y}$ including other tokens beyond the answer, e.g., $\mathbf{y}=(\mathrm{\_185},\mathrm{6},\mathrm{\_minus},\mathrm{\_2},\mathrm{\_equals},\mathrm{\_185},\mathrm{4})$. In both cases, the $\mathrm{extract}$ function must parse and detokenize the output; in CoT, there is some extra work to spot where the answer is placed.
Our prompts can explicitly encourage use of direct answer or chain of thought as strategies, which we denote as $\mathcal{I}_{\mathrm{da}}$ and $\mathcal{I}_{\mathrm{cot}}$. For eliciting CoT, this includes strategies like telling a model to “think step by step(Kojima et al., 2022). For directly answering a question, a prompt may say “immediately generate the answer”. We track the average location of the answer in the generated output for both CoT and direct prompts in Appendix F.3 to ensure that direct answer prompts give the answer early in the output. We also ensure that $\mathrm{extract}$ can parse answers from the generated output for each model, prompt, and dataset combination used in our experiments, tailoring the $\mathrm{extract}$ function as needed to ensure low unparseable rates for each model and task. [^3] We exclude a number of other “CoT-like” approaches in our analysis such as decomposed prompting (Khot et al., 2023; Zheng et al., 2024a) and multi-agent debate (Du et al., 2023; Chen et al., 2024). We focus on single prompt approaches. We deal with tool-augmented approaches in Section 5. All prompts and outputs per dataset per model have been uploaded to Huggingface and we include examples of some of our prompts in the Appendix J. We also experiment with few-shot CoT prompts, which we find perform similarly to zero-shot prompts; details about these are given in Appendix E.
@@
Central claim
Figure 1 shows that there are a large number of positive results on CoT reported in the literature. Informally, we believe many readers of the literature to hold the following view: $\mathcal{I}_{\mathrm{cot}}$ will outperform $\mathcal{I}_{\mathrm{da}}$ on nearly all reasoning problems, whether those problems involve symbolic or non-symbolic reasoning. Our evidence does not support this conjecture. We will show that this performance boost is strongest for symbolic and semi-symbolic tasks, while giving little to no improvement (or even hurting performance) on non-symbolic tasks.
@@ 3 Results from the Literature
Criteria and Process
We investigate all papers from ICLR 2024, a representative ML venue, and two representative NLP venues, EACL 2024 and NAACL 2024 (including Findings and Workshop papers). This resulted in 4,642 papers total that filtered using automatic and manual methods to papers including experiments comparing chain-of-thought, $\mathcal{I}_{\mathrm{cot}}$, against direct answering prompts, $\mathcal{I}_{\mathrm{direct}}$. A total of 110 papers were found that matched our criteria with 1,218 experimental comparisons. We then grouped the comparisons by the types of tasks and datasets being evaluated. More details on our automatic and manual filtering, as well as our categorization, can be found in Appendix A and B.
Table 1: A few categories for experimental comparisons. Full list in Appendix B.
Category Description
Logical reasoning Tasks designed to test for logical reasoning, whether deductive (Saparov & He, 2023, PrOntoQA), inductive (Bowen et al., 2024) or analogical (Ma et al., 2024) reasoning, including syllogisms and logical puzzles.
Results
Do any non-math datasets benefit from CoT?
There are a few outliers that less clearly follow the trend. ScienceQA (Lu et al., 2022) consists of multiple choice questions across a range of natural and social science disciplines, though it is hard to interpret gains without knowing breaking down performance by subject or question type. The dialogue evaluation dataset from Jia et al. (2024) sees large improvements with CoT, but this is a proprietary dataset, and we note that other essay scoring results in our meta-analysis (Li et al., 2024; Stahl et al., 2024) did not show improvements with CoT. Other non-math, symbolic or logical datasets that benefit from CoT are Commitment Bank (de Marneffe et al., 2019) and the task of eliciting verbalized confidence (Xiong et al., 2024). Nevertheless, these are exceptions to the rule. The majority of the reported benefits from using CoT in the NLP and ML literature comes from math or math-related tasks.
4 Results from Experiments
4.1 Experimental Setup
Dataset, Models, Prompts
All datasets, models, and prompts we evaluate over can be found in detail in the tables 3, 4, and 5 of Appendix C. We restricted our experiments to English models commonly used and benchmarked on general reasoning datasets. Our datasets include those which are widely used in CoT and reasoning literature, including a mix of non-symbolic, semisymbolic, and symbolic reasoning. They span different formats, including multiple-choice, short-answer, and free-response; however, most of these datasets are multiple choice or short answer, as CoT is not typically used in long-form response settings. We also categorize each dataset into a larger category of reasoning required to solve it: Commonsense, Knowledge, Symbolic, Mathematical, and Soft Reasoning. We define Soft Reasoning as questions relying on commonsense and natural language but going beyond simple inferences about these statements. Finally, we explore several prompting strategies for eliciting reasoning from language models, as past work has emphasized the importance of the prompt (Yang et al., 2024). However, we generally found slight performance differences; see Appendix D for details. We therefore focus on prompts similar to Kojima et al. (2022) and Wei et al. (2022) for zero-shot and few-shot settings, respectively, with alterations to improve the model’s ability to produce desired behavior (i.e., formats that allow for easily parsed answers). We upload all our prompts and outputs for each model for each prompting strategy on Huggingface [^4] https://huggingface.co/collections/TAUR-Lab/cot-analysis-project-66bbb9e5e0156e65059895f5.
Implementation Details
We use a high-throughput inference package, vLLM (Kwon et al., 2023), for the model inference process. We use greedy decoding on all models. Our prompts are taken from the Llama 3.1 evaluations when available (Dubey et al., 2024), and minor adjustments are made to unify prompting strategies. For other datasets, we either use the standard prompt for the dataset from the corresponding original paper or implement our own prompt. Our answer parser ($\mathrm{extract}$) is tailored to each dataset and model. Specific details about each dataset, its prompts, and answer extractor can be found in Appendix C.
Figure 3: Left: Performance gain from using CoT for each reasoning category. Right: Performance gain from using CoT for each dataset, averaged across models and broken out across 5 representative models. Red lines indicate median improvement. In both plots we see a consistent trend: most improvements from using CoT are from math and symbolic reasoning.
4.2 Results
Where does zero-shot CoT improve over direct prompts?
By contrast, the mathematical and symbolic categories get larger boosts in improvements alongside symbolic and many semi-symbolic datasets. MATH and GSM8K show gains as large as 41.6% and 66.9%, respectively. The semi-symbolic datasets like ContextHub and MuSR Murder Mysteries show moderate gains. These datasets require the application of logical rules to reach the answer, e.g., first-order logic parsed from simple natural language (ContextHub) or more complex commonsense statements (MuSR Murder Mysteries). All results are shown in the Appendix F.1 as well as a full list of numeric results for both CoT and direct answer prompting in Table LABEL:tab:raw_exp_0_zs_results. We also explored the few-shot setting and found it had little impact on when CoT will help; see Appendix E.
Does the answer format impact where CoT will help?
Many of the commonly-used datasets for problems other than math are multiple choice. We highlight here that CoT has similar performance to direct answer across models for two datasets that are not multiple-choice and contain varying levels of non-symbolic reasoning. First, MuSiQue (Trivedi et al., 2022) is a short-form QA task requiring multi-hop reasoning. We consider this a semi-symbolic dataset as the questions have an explicit multi-hop structure. Because answer spans in MuSiQue can be paraphrased in many different ways, we use GPT-4o to judge if two answer spans are equivalent. Despite being semi-symbolic, we see no overall improvement from CoT.
Mostly no, except for MMLU, StrategyQA, and MuSR.
We tested the significance of the improvements from CoT on the 13 datasets in the Knowledge, Soft Reasoning, and Commonsense reasoning categories using paired bootstrapping to assess whether CoT gives a significant improvement. To account for multiple comparisons, we applied a Bonferroni correction, setting the p-value to $0.00027$ to account for the 14 models and 13 datasets. About 32% (59) of the datasets that show a benefit in these three reasoning categories were considered significant. Nearly half of these comparisons (26) are on MMLU and MMLU Pro. On these datasets, we find that CoT is mainly helping on math-related questions. StrategyQA and MuSR also received a consistent performance boost across 10 and 6 models respectively. StrategyQA is often used to benchmark reasoning methods and is built specifically to get a benefit from methods that decompose the question into steps, so a gain in performance is not unprecedented. MuSR, similarly, was built to have multiple steps of complex natural language reasoning, which may receive benefits from CoT. The remaining datasets that receive significant benefits are spread across the datasets and models.
Why do MMLU and MMLU Pro get a boost?
MMLU and MMLU Pro contain many different questions requiring different types of reasoning. We separated MMLU and MMLU Pro questions into two bins, those related to math and those not related to math, by checking if the questions text or generated response from the LLM includes an “=”. Figure 4 shows that a majority of the performance gain seen from MMLU and MMLU Pro is from the math slices of each dataset. See more details in Appendix G.
5 Strengths and weaknesses of CoT at formal reasoning
By separating planning and execution in this way, we can test how much a language model can gain from only having a plan, to having a plan and solving it with CoT, or to having a plan and then solving it with an external symbolic solver. Given a plan $\mathcal{S}_{\mathrm{plan}}\sim\mathcal{I}^{m}_{\mathrm{planning}}(\mathbf{q})$, we compare the performance of the settings below to evaluate at which stage LM is most effective and falls short.
Figure 5: Prompt variants that separate planning and execution for GSM8K. For all prompt variants besides direct answer and CoT (not shown), we few-shot prompt an LLM to first generate a Python program as a solution plan. For Plan + Direct Solver, the LLM is prompted to directly give an answer from the plan; for Plan + CoT Solver, the LLM is prompted to solve the plan step-by-step with CoT and give an answer; for Plan + Tool Solver, we feed the plan into a Python interpreter.
5.1 Settings Evaluated
Settings 1 and 2: Few-shot direct answer and CoT: We use the few-shot direct answer and CoT prompts from Section 4.1 as baselines. Figure 5 includes an example of each setting on GSM8K.
Settings 3 and 4: Plan + Direct Solver and Plan + CoT Solver: Here we use inspiration from Xu et al. (2024a) and generate a symbolic plan using the same strategy as Ye et al. (2023). Specifically, we use a few-shot prompt $\mathcal{I}^{m}_{\mathrm{planning}}$ to generate a formal specification $\mathcal{S}_{\mathrm{plan}}$ that should be executable by a symbolic solver. In the same prompt LMs are asked to solve their generated specification $\mathcal{S}_{\mathrm{plan}}$ and derive the final answer $\tilde{\mathbf{y}}\sim p(\mathbf{y}\mid\mathcal{I}_{\mathrm{da}}(\mathcal{S}_{\mathrm{plan}}))$, either directly giving the answer after generating the specification (Plan + Direct Solver) or providing step-by-step explanations and tracking of intermediate steps for the derivation (Plan + CoT Solver). Particularly, $\mathcal{S}_{\mathrm{plan}}$ is a Python program for math datasets, and is a set of first-order logic specifications for logical reasoning datasets.
Setting 5: Plan + Tool Solver: We then evaluate how effective CoT can be at performing symbolic computations compared with external symbolic solvers. Following prior work on augmenting LMs with tools for math and logic questions (Ye et al., 2023; Pan et al., 2023; Gao et al., 2023; Chen et al., 2023), we generate $\mathcal{S}_{\mathrm{plan}}$ the same way as in CoT Solver, but now feed in the plan into a symbolic solver (Python interpreter or a SMT Solver), such that $\hat{a}=\mathrm{solve}(\mathcal{S}_{\mathrm{plan}})$.
Evaluation Setup: We compare the performance of each setting on math (GSM8K) and logical reasoning (ContextHub and FOLIO) datasets. We follow Gao et al. (2023) to include GSM8K-Hard, a minimally modified version that replaces numbers of GSM8K with larger numbers, to account for the possibility of recent LLMs overfitting GSM8K by data contamination (Zhang et al., 2024).
For Plan + Direct solver and Plan + CoT solver, we use the few-shot prompts from Ye et al. (2023). For Plan + Tool solver, we use state-of-the-art tool-augmented prompting methods. Particularly, for GSM8K, we use Program-aided Language Model (Gao et al., 2023, PAL) that executes the LM-generated plan with a Python interpreter. For logical reasoning datasets, we use Satisfiability-Aided Language Model (Ye et al., 2023, SatLM) that uses automated theorem prover Z3 (De Moura & Bjørner, 2008) to solve the generated specifications. If the generated plan cannot be parsed by the tool, we use random guessing when the question is multiple choice, and mark it incorrect otherwise.
5.2 Evaluation Results
Figure 6 shows the results across a representative selection of models. Detailed numerical results, including the unparseable rates of model-generated plans, can be found in Appendix H.
We argue that these results provide an explanation of why CoT helps on symbolic tasks. While all tasks could feasibly benefit from a detailed description of how to solve each individual question (e.g., a plan in the context of this section), CoT only outperforms direct answer when these steps require a substantial amount of tracing and computation. In these settings, we can see clear performance benefit from using symbolic solvers; CoT appears to be a poor (but universal) approximation to such solvers. When possible, LLMs should be paired with symbolic solvers at inference time when solving symbolic tasks to achieve consistently better performance over direct answer and CoT.
6 Discussion and Related Work
Where is CoT helping and why?
Can we improve CoT further?
Our work treats chain-of-thought variants that explicitly don’t involve multiple inferences. There is evidence that using additional calls to LLMs can help (Du et al., 2023; Yao et al., 2023; Besta et al., 2023; Chen et al., 2024), but these methods use significantly increased computation, and careful benchmarking sometimes reveals that naive techniques are as good as iterative ones (Olausson et al., 2024). However, past theoretical results show that Transformers are augmented in a fundamental way by CoT (Liu et al., 2023b; Merrill & Sabharwal, 2024); we believe this indicates the potential for improving CoT beyond prompt-based CoT. On the other hand, recent methods showing benefit from “internalizing” CoT (Deng et al., 2024) may indicate that explicit generation of intermediate tokens is not used to its full potential.
Limitations
One set of tasks we do not cover in our experiments (except for BiGGen Bench) is long-horizon planning. However, many works in the literature have already discussed the efficacy of planning with CoT. We also do not address the data contamination of some of these models on the datasets. We try to mitigate this by including multiple models, datasets (new and old), and our meta-analysis. For more discussion of planning and dataset contamination, see Appendix I.
7 Conclusion
In this work, we characterize the performance of prompt-based CoT through a meta-analysis of the literature and experiments across different models, datasets, and prompts. We find that CoT predominantly helps on math and formal logic, largely due to its ability to trace the intermediate steps of a problem. But CoT rarely outperforms tool-augmented approaches for these same problems. We believe that CoT remains a powerful technique, but to give improvement across a wider range of NLP tasks, research should move beyond prompt-based CoT to new paradigms like search, interacting agents, or better fine-tuned models.
Reproducibility
For our experiments, we provide in-depth details of how we evaluated models on each dataset in Section 4.1 and Appendix C. Furthermore, we release all prompts for every dataset on Huggingface, including per model output and sampling parameters. For our meta-analysis of the literature, we describe our filtering criteria and process of annotating experiments into high-level categories in Section 3 and Appendix B. We also release the full list of papers in our meta-analysis together with extracted experimental comparisons and task category annotations.
Appendix A Meta-analysis Expanded Details on Criteria and Process
@@
This resulted in a total of 1,218 experimental comparisons across 110 papers (35 from ICLR and 75 from NAACL and EACL) covering 264 datasets. Details and more information will be available in our GitHub Repo.
Categorization
Given the large number of tasks and datasets being compared, we grouped each task into a set of 14 categories. These categories were determined based on the description (and possibly examples) of the task, not taking into account system performance. These categories abstract over traditional NLP task classifications (e.g., NER, reading comprehension) and take into account both the task format and the kinds of reasoning involved. Definitions for several categories are shown in Table 1 and the full description is given in Appendix B.
Appendix B Quantitative Meta-Analysis
@@
Logical reasoning Tasks designed to test for logical reasoning, whether deductive (Saparov & He, 2023, PrOntoQA), inductive (Bowen et al., 2024) or analogical (Ma et al., 2024) reasoning, including syllogisms and logical puzzles.
@@
Table 2: Categories and their descriptions for the meta-analysis.
Appendix C Expanded Experimental Details